I am brain dead today and could use help here please! Thanks so much for any help you can give.
A student makes a $9.25 purchase at a bookstore with a $20. bill. The store has no bills and gives the change in quarters and fifty cent pieces. There are 30 coins in all. How many of each kind are there?

Answer by scott8148(5895)

let x=quarters, so 30-x="fifty cent pieces" __ working in cents to avoid decimals

(25)(x)+(50)(30-x)=2000-925 __ 25x+1500-50x=1075 __ -25x=-425